The Destruction of Jericho The Bible describes the Exodus

as a period of 40 years that began with plagues in Egypt

and ended with the destruction of Jericho. Archeologists

seeking to establish the exact dates of these events have

proposed that the plagues coincided with a huge eruption

of the volcano Thera in the Aegean Sea.

a. Radiocarbon dating suggests that the eruption occurred

around 1360 bce, though other records place the

eruption of Thera in the year 1628 bce. What is the

percent difference in the 14C decay rate in biological

samples from these two dates?



1
Expert's answer
2021-05-28T00:58:51-0400

k = ln2 / half-life.


The half life of carbon is 5730 years.


So k = ln2/5730 yr = 1.2097 × 10-4 yr-1


Now, substitute the values the formula. we will get


14C​2007AD/14C​1315 BC = e-(1.2097 ×10-4)(3322)


14C​2007AD/14C​1315 BC = 0.6690728


So the ratio of 14C in comparison of 2007 AD to that of 1315 BC = 0.6690728